# Sharpness-aware Parameter Selection for Machine Unlearning

### v1

@@ 2 Related work
Another strain of machine unlearning is approximate unlearning, which usually achieves unlearning by conducting several gradient updates. The most straightforward approach is gradient ascent over target data Golatkar et al. (2020). Several works discuss efficiency of the approximate unlearning. Jia et al. (2023) showed that conventional model sparsification strategies, such as weight pruning, improve machine unlearning. Specifically, the authors demonstrated that a model unlearned after pruning exhibited smaller performance gap with the retrained model. Along with this finding, the authors proposed machine unlearning frameworks aided by sparsification. Recently, Li et al. (2025) proposed an efficient unlearning by reducing the number of data samples to unlearn. The authors identified that certain samples do not fundamentally change the model behavior after unlearning. Thus, by filtering out such unnecessary data, the authors are able to efficiently achieve unlearning by only removing crucial data samples.
@@
For large language models, increasing efficiency of unlearning frameworks is even more crucial due to their large number of parameters. Similar to fine-tuning, LoRA (Hu et al., 2022) is widely used to unlearning large language models. Gao et al. (2024) proposed a practical unlearning algorithm for large language models. It freezes the original weights and fine-tunes LoRA with target prompts to deter the model from providing the original response. Then, it utilizes a classification model to determine the similarity of given prompts with the unlearning contents. Depending on the similarity, adapters are activated to emit the safe answers. Similarly, Cha et al. (2024) utilized gradient ascent with LoRA. Instead of directly utilizing gradient ascent on LoRA, the authors further proposed additional hinge loss to boost unlearning stability. Hu et al. (2024b) proposed exact unlearning on LLMs. Instead of training a separate model on each data partition, the authors also partition adapters and train each adapter partition on each corresponding data partition. Recently, Jia et al. (2025) proposed a parameter selection strategy for large language model unlearning. The authors score the model weights by their gradient sensitivity. Specifically, they introduce small changes to the gradients, then select the weights that demonstrate significant response to the change, which are considered critical for unlearning. However, this requires solving complex optimization problems, which could lower the overall efficiency.
@@
Hessians, or the second order gradients are widely utilized on machine unlearning. Primarily, certified unlearning Guo et al. (2019) requires the inverse of Hessians to analyze the gradient norm and bounding unlearning budget, which is expensive to compute regardless of the complexity of the model Zhang et al. (2024a). Some works utilized Hessian approximations to increase efficiency Mehta et al. (2022). Unlike our work, the primarily focus of Hessians in certified unlearning is to complete the mathematical analysis, rather than finding the salient parameters. Hessians are rarely utilized for salient parameter selection. Liu et al. (2023) used the fisher information matrix, or expected hessian matrix to selectively choose parameters to update. The authors utilize influence function and derive upper bound of Kullback-Leibler divergence between original parameters and estimated retrained parameters. The upper bound is determined by diagonals of the fisher information matrix. The authors choose parameters that can lower the upper bound. However, although the paper empirically show their approach is effective for deep models, the upper bound analysis is only valid to linear models. Some works conduct gradient ascent assisted by inverse Hessian matrix, to accelerate optimization of parameters containing knowledge of unlearning samples from large language models Jia et al. (2024); Gu et al. (2024). Most of them utilized the hessian matrix in the broad context of second order update. While few works choose parameters via fisher information matrix, and some attempted parameter selection via first order gradient Fan et al. (2023). Recently,  Foster et al. (2024) proposed selective synaptic dampening (SSD), a salient parameter selection using fisher information matrix. The authors obtain two Fisher information matrices: one from the retain data and one from the forget data. They then select the parameters whose corresponding elements in the Fisher information matrix from the forgotten data are higher than those in the matrix from the retained data. Huang et al. (2024) proposed unlearning with steepest descent and also used the same selection strategy. While this is closely related to our work, it is not directly applicable to our problem settings. SSD expects the sample unlearning scenario (unlearning both features and labels) while our problem setting is the feature unlearning (unlearn a subset of features or labels).
@@ 4 Sharpness-Aware Parameter Selection for Unlearning
$\bm{\theta}^{u}$ $\approx\bm{\theta}^{*}-H_{\bm{\theta}^{*}}^{-1}\big{(}\nabla_{\bm{\theta}}L_{forget}(\bm{\theta}^{*};\mathcal{D}_{f},\tilde{\mathcal{D}_{f}})\big{)},$ (12)
where $H_{\bm{\theta}^{*}}$ is the Hessian of $L_{learn}(\bm{\theta};\mathcal{D}_{f},\tilde{\mathcal{D}_{f}})$ at $\bm{\theta}^{*}$. Therefore the above second-order update is not a Newton update, as the Hessian and the gradient vector are computed on two different losses $L_{learn}$ and $L_{forget}$, respectively. From the above update rule, we can clearly observe that those parameters in $\bm{\theta}^{*}$ which are located in a wide dip of $L_{learn}$ will experience a larger learning rate through the update process. Therefore, those are the parameters which will be updated the most. Therefore, our proposed parameter selection provides a valid approximation to the above closed-form second-order update for unlearning, by eliminating the need to compute the inverse of the complete Hessian matrix. In fact, our strategy needs only the diagonal elements of the Hessian matrix for performing its parameter selection, which can be computed with computational complexity of $\mathcal{O}(p)$.

### v2

@@ 2 Related work
Another strain of machine unlearning is approximate unlearning, which usually achieves unlearning by conducting several gradient updates. The most straightforward approach is gradient ascent over target data (Golatkar et al., 2020). Several works discuss efficiency of the approximate unlearning. Jia et al. (2023) showed that conventional model sparsification strategies, such as weight pruning, improve machine unlearning. Specifically, the authors demonstrated that a model unlearned after pruning exhibited smaller performance gap with the retrained model. Along with this finding, the authors proposed machine unlearning frameworks aided by sparsification. Recently, Li et al. (2025) proposed an efficient unlearning by reducing the number of data samples to unlearn. The authors identified that certain samples do not fundamentally change the model behavior after unlearning. Thus, by filtering out such unnecessary data, the authors are able to efficiently achieve unlearning by only removing crucial data samples.
@@
For large language models, increasing efficiency of unlearning frameworks is even more crucial due to their large number of parameters. Similar to fine-tuning, LoRA (Hu et al., 2022) is widely used for unlearning large language models. Gao et al. (2024) proposed a practical unlearning algorithm for large language models. It freezes the original weights and fine-tunes LoRA with target prompts to deter the model from providing the original response. Then, it utilizes a classification model to determine the similarity of given prompts with the unlearning contents. Depending on the similarity, adapters are activated to emit the safe answers. Similarly, Cha et al. (2024) utilized gradient ascent with LoRA. Instead of directly utilizing gradient ascent on LoRA, the authors further proposed additional hinge loss to boost unlearning stability. Hu et al. (2024b) proposed exact unlearning on LLMs. Instead of training a separate model on each data partition, the authors also partition adapters and train each adapter partition on each corresponding data partition. Recently, Jia et al. (2025) proposed a parameter selection strategy for large language model unlearning. The authors score the model weights by their gradient sensitivity. Specifically, they introduce small changes to the gradients, then select the weights that demonstrate significant response to the change, which are considered critical for unlearning. However, this requires solving complex optimization problems, which could lower the overall efficiency.
@@
Hessians, or the second order gradients are widely utilized in machine unlearning. Primarily, certified unlearning (Guo et al., 2019) requires the inverse of Hessians to analyze the gradient norm and bounding unlearning budget, which is expensive to compute regardless of the complexity of the model (Zhang et al., 2024a). Some works utilized Hessian approximations to increase efficiency (Mehta et al., 2022). Unlike our work, the primarily focus of Hessians in certified unlearning is to complete the mathematical analysis, rather than finding the salient parameters. Hessians are rarely utilized for salient parameter selection. Liu et al. (2023) used the fisher information matrix, or expected hessian matrix to selectively choose parameters to update. The authors utilize influence function and derive upper bound of Kullback-Leibler divergence between original parameters and estimated retrained parameters. The upper bound is determined by diagonals of the fisher information matrix. The authors choose parameters that can lower the upper bound. However, although the paper empirically show their approach is effective for deep models, the upper bound analysis is only valid to linear models. Some works conduct gradient ascent assisted by inverse Hessian matrix, to accelerate optimization of parameters containing knowledge of unlearning samples from large language models (Jia et al., 2024; Gu et al., 2024). Most of them utilized the hessian matrix in the broad context of second order update. While few works choose parameters via fisher information matrix, and some attempted parameter selection via first order gradient (Fan et al., 2023). Recently,  Foster et al. (2024) proposed selective synaptic dampening (SSD), a salient parameter selection using fisher information matrix. The authors obtain two Fisher information matrices: one from the retain data and one from the forget data. They then select the parameters whose corresponding elements in the Fisher information matrix from the forgotten data are higher than those in the matrix from the retained data. Huang et al. (2024) proposed unlearning with steepest descent and also used the same selection strategy. While this is closely related to our work, it is not directly applicable to our problem settings. SSD expects the sample unlearning scenario (unlearning both features and labels) while our problem setting is the feature unlearning (unlearn a subset of features or labels).
@@ 4 Sharpness-Aware Parameter Selection for Unlearning
$\bm{\theta}^{u}$ $\approx\bm{\theta}^{*}-H_{learn}^{-1}(\bm{\theta}^{*})\big{(}\nabla_{\bm{\theta}}L_{forget}(\bm{\theta}^{*};\mathcal{D}_{f},\tilde{\mathcal{D}_{f}})\big{)},$ (12)
where $H_{learn}(\bm{\theta}^{*})$ is the Hessian of $L_{learn}(\bm{\theta};\mathcal{D}_{r},\mathcal{D}_{f})$ at $\bm{\theta}^{*}$. Therefore the above second-order update is not a Newton update, as the Hessian and the gradient vector are computed on two different losses $L_{learn}$ and $L_{forget}$, respectively. From the above update rule, we can clearly observe that those parameters in $\bm{\theta}^{*}$ which are located in a wide dip of $L_{learn}$ will experience a larger learning rate through the update process. Therefore, those are the parameters which will be updated the most. Therefore, our proposed parameter selection provides a valid approximation to the above closed-form second-order update for unlearning, by eliminating the need to compute the inverse of the complete Hessian matrix. In fact, our strategy needs only the diagonal elements of the Hessian matrix for performing its parameter selection, which can be computed with computational complexity of $\mathcal{O}(p)$.